1. A 35 πœ‡πΆ point charge is placed 32 π‘π‘š from an identical 32 πœ‡πΆ charge. How much work would be required to move a 0.50 πœ‡πΆ charge from a point midway between them to a point 12 π‘π‘š closer to either of the charges?

2. Three charges are at the vertices of an isosceles triangle. With π‘ž=7.00 𝑛𝐢, the two charges at the base of the triangle are separated by 2.00 π‘π‘š and each have a charge of βˆ’π‘ž. The charge q at the top of the triangle is 4.00 π‘π‘š from each of the charges at the base. Calculate the electric potential halfway between the charges on the base of the triangle.

3. (a) Show that the capacitance of a parallel-plate capacitor is given by the equation

C=EA/D

(b) A 1-megabit computer memory chip contains many 60.0Γ—10βˆ’15 𝐹 capacitors. Each capacitor has a plate area of 21.0Γ—10βˆ’12 π‘š2. Determine the plate separation of such a capacitor. (Assume a parallel-plate configuration.) The diameter of atom is on the order of 10βˆ’10 π‘š = 1 ˚ 𝐴. Express the plate separation in angstroms.


1
Expert's answer
2021-04-29T10:41:52-0400

(1) "q_1=q_2=q=35\mu C"

"q_3=0.50\mu C"

Work Done "=qV=q_3(V_b-V_a)"

"(0.50\mu C)(k\times35\mu C)\Bigg[\left(\dfrac{1}{0.04\space m}+\dfrac{1}{0.28\space m}\right)-\left(\dfrac{1}{0.16\space m}+\dfrac{1}{0.16\space m}\right)\Bigg]"

"=2.53\space J"


(2) "V_1=\dfrac{kq}{r_1}=\dfrac{9\times10^9\times7\times10^{-9}}{\sqrt{15}\times10^{-2}}=1626.65\space V"

"V_2=\dfrac{k(-q)}{r_2}=\dfrac{9\times10^9\times(-7)\times10^{-9}}{{1}\times10^{-2}}=-6300\space V"

"V_3=\dfrac{k(-q)}{r_3}=\dfrac{9\times10^9\times(-7)\times10^{-9}}{{1}\times10^{-2}}=-6300\space V"

"\sum V=V_1+V_2+V_3=-10973.35\space V"


(3)


The flux through the curved part outside the plates is also zero as the direction of the field E is parallel to this surface. The flux through "\Delta A" is

"\Phi=\vec{E}.\Delta \vec A=E\Delta A"

The only charge inside the Gaussian surface is

"\Delta Q=\sigma\Delta A=\dfrac{Q}{A}\Delta A"


From Gauss’s law,

"\oint\vec E.d\vec S=\dfrac{Q_{in}}{\epsilon_o}"

"E\Delta A=\dfrac{Q}{\epsilon_oA}\Delta A"

"E=\dfrac{Q}{\epsilon_o A}"


The potential difference between the plates is

"V=V_+-V_-=-\int_{A}^{B} \vec E \,.d\vec r"

"V=\int_A^BEdr"

"V=Ed=\dfrac{Qd}{\epsilon_o A}"


The capacitance of the parallel-plate capacitor is

"C=\dfrac{Q}{V}=\dfrac{Q\epsilon_o A}{Qd}"

"C=\dfrac{\epsilon_o A}{d}"


(b) "C=60\times10^{-15}\space F"

Area, "A=21\times10^{-12}\space m^2"

"d=\dfrac{\epsilon_o A}{C}=\dfrac{8.85\times10^{-12}\times21\times10^{-12}}{60\times10^{-15}}"

"d=3.0975\times10^{-9}\space m=30.975\space A^o"